A random variable X can assume two values 1 and 3. If its mean is 8/3, compute: the probabilities for those two points and the variance of X
Answer by stanbon(57361)

A random variable X can assume two values 1 and 3. If its mean is 8/3, compute: the probabilities for those two points
----------------
Let probability of "1" be x
Then probability of "3" is 1-x
----------------------------------
Mean = 1*x + 3(1-x) = 8/3
x + 3 - 3x = 8/3
-2x = 8/3 - 9/3
-2x = -1/3
x = 1/6 (Probability of "1")
1-x = 5/6 (Probability of "3")
==================================
I'll let you figure the variance.
==================================
Cheers,
Stan H.